OpenStudy (tdskitz):

Solve for x: |2x + 6| − 4 = 20 A| x = 9 and x = 11 B| x = −9 and x = 15 C| x = 9 and x = −15 D| No solution

OpenStudy (google):

Part 1: 2x+6-4=20 2x+2=20 2x=20-2 x=18/2 x=9

OpenStudy (google):

Part 2: -(2x+6)-4=20 -2x-6-4=20 -2x-10=20 -2x=20+10 -2x=30 x=-15

OpenStudy (google):

What would your answer be?

OpenStudy (tdskitz):

C?

OpenStudy (google):

Correct!